Tube B42 has a concentration 2 x 106/mL of HeLa cells. A 24-well plate is to be seeded with 1 x 105 cells per well. Your experiment uses 500μL total volume (cells + DMEM media) per well in the plate. • What volumes of tube B42 and DMEM of need to be added to each well?
Answer #1

the concentration of cells is 2*10^6 cells/mL

and in each well final volume= 500uL

M1= 2*10^6 cells/mL

V2= 500uL

M2= 1*10^5 cells/mL

V1= ?

M1V1=M2V2

V1= M2V2/M1

= 1*10^5*500/2*10^6

= 50 uL

so take 50uL of B42 cells stock in the well and add ( 500-50=450 mL ) 450 mL of DMEM to each well.